### v1

# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

Abstract
Fine-tuning large pretrained models on private datasets may run the risk of violating privacy. Differential privacy is a framework for mitigating privacy risks by enforcing algorithmic stability. DP-SGD enables training models with private data in a privacy-preserving manner, but raises new obstacles in the form of performance loss and significant engineering challenges. We introduce DP-ZO, a new method for fine-tuning large language models that preserves the privacy of training data by privatizing zeroth-order optimization. A key insight into the design of our method is that the direction of the gradient in SPSA, the zeroth-order algorithm we use, is always random and the only information that depends on private data is the step size, i.e., a scalar. Therefore, we only need to privatize the scalar step size, which is memory-efficient. DP-ZO, which can be instantiated with either Laplace or Gaussian noise, provides a strong privacy-utility trade-off across different tasks, and model sizes, under conservative privacy budgets. One noteworthy result is that DP-ZO exhibits just $1.86\%$ performance degradation due to privacy at $(1,10^{-5})$-DP when fine-tuning OPT-66B on 1000 training samples from SQuAD.
1 INTRODUCTION
The proliferation of open-source models pretrained on web-scale datasets (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023) has created a paradigm shift in privacy preserving machine learning. Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for preserving privacy while training models on private data, but it requires additional data (Tramèr and Boneh, 2021) to prevent a drop in utility (Yu et al., 2021a). Pretrained model checkpoints have emerged as a compelling “free” source of prior information to boost the performance of DP training (Ganesh et al., 2023; Tang et al., 2023a; Panda et al., 2022). By only requiring DP during the fine-tuning phase, a recent line of work (Li et al., 2022b, a; Yu et al., 2021b; He et al., 2023; Bu et al., 2023c) is able to obtain impressive performance with DP-SGD (Abadi et al., 2016). Despite these advancements, DP needs additional engineering effort and may not be available in real-world deployments of fine-tuning APIs (OpenAI, 2023; Anyscale, 2023). We propose a new direction for DP fine-tuning of large pretrained models that achieves strong privacy-utility trade-off and is more resource-efficient, easy to implement, and portable.
In this work, we introduce a new methodology for DP fine-tuning of large pretrained models. Our method uses zeroth-order optimization (ZO) (Spall, 1992). A key insight into this ZO optimization method is the power of two forward passes, each done with the same random perturbation with flipped signs. The difference in losses between the two perturbations tells us which perturbation is better, and we update the model by taking a step in that direction. This approach, which we describe in detail in Section 2.2, allows us to approximate the gradient without backpropagation.
Our differentially private zeroth order (DP-ZO) method privatizes the zeroth-order update, by adding noise to the difference between the losses. This noise is proportional to the sensitivity of this loss difference with respect to changing a single example in the training set, which is controlled by clipping. We limit the $\ell_{p}$ sensitivity by clipping the norm of the difference in scalar losses, between the two random perturbations. By removing the need for per-example gradient clipping (Abadi et al., 2016), our method enables DP training of large models with just a few lines of code.
[…]
2 BACKGROUND
2.1 Differential Privacy
Differential privacy (DP) is the gold standard method for providing algorithmic privacy (Dwork et al., 2006).
Definition 2.1 ($(\varepsilon,\delta)-$ Differential Privacy (DP)).
An algorithm $\mathcal{M}$ is said to be $(\varepsilon,\delta)$-DP if for all sets of events $S\subseteq\text{Range}(\mathcal{M})$ and neighboring datasets $D,D^{\prime}\in\mathcal{D}^{n}$ (where $\mathcal{D}$ is the set of all possible data points) we have the guarantee:
$\Pr[\mathcal{M}(D)\in S]\leq e^{\varepsilon}\Pr[\mathcal{M}(D^{\prime})\in S]+\delta$ (1)
where $D$ and $D^{\prime}$ are neighboring datasets if they differ in a single entry, $\varepsilon$ is the privacy budget, and $\delta$ is the failure probability. When $\delta=0$, we term it as pure $(\varepsilon,0)$-DP or $\varepsilon$-DP for simplicity.
We define a set of existing differential private mechanisms that we will use in our work.
Proposition 2.2 (Gaussian mechanism (Dwork and Roth, 2014)).
For any function $f:\mathbb{X}^{n}\rightarrow\mathbb{R}$ with $l_{2}$ sensitivity $\Delta$, the mechanism defined as
$M(X)=f(X)+z,$
where $z\sim N\left(0,\frac{2\ln(1.25/\delta)\Delta^{2}}{\varepsilon^{2}}\right)$, provides $(\varepsilon,\delta)$-DP.
Proposition 2.3 (Laplace mechanism (Dwork and Roth, 2014)).
For any function $f:\mathbb{X}^{n}\rightarrow\mathbb{R}$ with $l_{1}$ sensitivity $\Delta$ the mechanism defined as
$M(X)=f(X)+z,$
[…]
2.2 Zeroth-order Optimization
We use a method from the large body of work on zeroth-order optimization (Flaxman et al., 2004; Shamir, 2013; Ghadimi and Lan, 2013; Nesterov and Spokoiny, 2017) that uses the difference in losses between two random perturbations (Duchi et al., 2015; Spall, 1992) with opposite signs to determine the magnitude of a gradient update in the direction of the random perturbations. In the non-private setting where the adaptation between the pretrained model and the fine-tuning dataset has low rank (Hu et al., 2022), as in fine-tuning large language models, Malladi et al. (2023) show this method converges at a rate that is not catastrophically slower than SGD fine-tuning.
Definition 2.4 (Simultaneous Perturbation Stochastic Approximation (SPSA)  (Spall, 1992))).
Given a model with parameters $\theta\in\mathbb{R}^{d}$ and a loss function $\mathcal{L}$, the gradient estimate on a minibatch drawn from a dataset $\mathcal{B}\subset\mathcal{D}$ is computed by projecting the loss on the minibatch $\mathcal{L}(\theta;\mathcal{B})$ onto a random perturbation $z\in\mathbb{R}^{d}$ that is a standard Gaussian random variable (i.e., $z\sim\mathcal{N}(0,I_{d})$) scaled by $\phi$:
$\hat{\nabla\mathcal{L}_{b}}(\theta;\mathcal{B})=$ $\frac{\mathcal{L}(\theta+\phi z;\mathcal{B})-\mathcal{L}(\theta-\phi z;\mathcal{B})}{2\phi}z$ (2)
Random perturbations in zeroth-order optimization (ZO) serve as high-variance estimates of the actual gradient, enabling optimization without the need for explicit gradient computations. However, these perturbations themselves carry a privacy risk. The characteristics of the perturbations can be inferred from the gradient updates, effectively leaking information about the data. Moreover, the magnitude of the perturbations can be estimated with high precision in high dimensions due to the central limit theorem. Therefore, incorporating differential privacy into ZO is essential to safeguard against these vulnerabilities.
3 OUR METHOD: DP-ZO
We introduce our framework for differentially private zeroth order optimization (DP-ZO) by integrating DP into Definition 2.4. In our framework, the information obtained from training data can be represented as a scalar. This scalar has a bounded sensitivity (when applying clipping) and can be privatized by adding noise. We expect this noise to be less detrimental to the utility, compared to the noise added in DP-SGD, as we are adding noise to a single dimension. In other words, we would expect the gap between non-private and private utility to be smaller than that of DP-SGD. However, it is possible that for some tasks the non-private performance of zeroth-order optimization is poor.
Algorithm 1 Differentially Private-ZO
1:Model parameters $\theta$, dataset $\mathcal{D}$, learning rate $\eta$, perturbation scale $\phi$, privacy parameter $\sigma$, noising mechanism $\mathcal{Z}$, clipping threshold ${C}$, expected batch size $B$, sub-sampling rate $p=B/|\mathcal{D}|$).
2:$g={0}$
3:for $t\in 1,\ldots T$ do
4:     Poisson sample $\mathcal{B}$ from $\mathcal{D}$ with sub-sampling rate $p$
5:     $\vec{z}\sim\mathcal{N}(\vec{0}_{|\theta|},\mathbf{I}_{|\theta|\times|\theta|})$
6:     $\theta^{+}\leftarrow\theta+\phi\vec{z}$
7:     $\theta^{-}\leftarrow\theta-\phi\vec{z}$
8:     for $(x_{i},y_{i})\in\mathcal{B}$ do
9:         $l_{i}^{+}\leftarrow\mathcal{L}(\theta^{+},(x_{i},y_{i}))$
10:         $l_{i}^{-}\leftarrow\mathcal{L}(\theta^{-},(x_{i},y_{i}))$
11:         $l_{i}=clip(l_{i}^{+}-l_{i}^{-},C)$
12:     $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+\mathcal{Z}(\mathrm{C},\sigma)}{B\cdot 2\phi}$
13:     $\theta=\theta-\eta s\vec{z}$
DP-ZO.
We explain the steps of our algorithm while emphasizing the key differences from Definition 2.4 required to guarantee $(\varepsilon,\delta)$-DP. We first sample a batch from the dataset with Poisson sampling (Balle et al., 2018) which allows us to use privacy amplification by subsampling. For each model parameter $\theta_{i}$ we want to update, we independently sample a perturbation $z_{i}$ from a standard Gaussian distribution and scale it by a predetermined constant $\phi$; we denote the full perturbation vector as $\phi\vec{z}$. Now we compute an approximation of the gradient by projecting it onto the random perturbation $\vec{z}$. That is, for a training sample $x_{i}$ we compute the difference in scalar losses between ${\theta+\phi\vec{z},\theta-\phi\vec{z}}$. Intuitively, this scalar tells us how much better one random step is than the other. We clip this scalar to limit the sensitivity (described in detail in the next paragraph). We add noise to the aggregation over samples in our training batch (described in detail in the subsequent paragraph). Finally, we take a step in the direction of $\vec{z}$ by scaling our private step size by the expected batch size, perturbation constant $\phi$, and the learning rate $\eta$.
DP-ZO enables new mechanisms by privatizing the difference in losses between perturbations.
DP-ZO proposes an update direction determined by a $d$-dimensional random vector, generated from a public PRNG, making the direction itself public information. The only private aspect is the step size, that is influenced by the difference in losses between perturbations with opposite signs. To privatize this step size, we add noise proportional to the sensitivity of the step size. We bound the sensitivity of the step size by clipping the per-example step sizes to a specific range $[-C,C]$, so the sensitivity under add-remove DP is $C$.
Given a private scalar with bounded sensitivity, we can apply the classical Gaussian mechanism to release a privatized scalar with $(\varepsilon,\delta)$-DP. The Gaussian mechanism is widely studied in privacy-preserving machine learning techniques like DP-SGD, in part because the best accounting techniques for the Gaussian mechanism (Dong et al., 2019; Gopi et al., 2021) are tight. However, the Gaussian mechanism can only provide $(\varepsilon,\delta)$-DP. To avoid such a failure case in extremely sensitive applications, researchers often recommend using cryptographically small values of $\delta$ (Vadhan, 2017). Unfortunately, due to limitations of accounting methods, we currently cannot calculate the tight privacy of composition of sub-sampled Gaussian mechanism for values of $\delta$ smaller than $10^{-10}$. Alternatively, we can resort to mechanisms that can obtain pure $\varepsilon$-DP. These mechanisms, such as Laplace mechanism, come with a guarantee that the mechanism will never fail catastrophically. However, due to large tails of the Laplace mechanism, it has never been a contender for high dimensional optimization.
Although it is possible to obtain pure DP with DP-SGD by adding Laplace noise scaled to the $\ell_{1}$ sensitivity of the gradient, this is challenging for large models because the $\ell_{1}$ sensitivity can be $\sqrt{d}$ times larger than the $\ell_{2}$ sensitivity (and often is), especially in the modern large machine learning models that consist of billions of parameters. In contrast, DP-ZO only requires to privatize the loss. The one-dimensional private estimation of the step size is amenable to the Laplace mechanism, because the $\ell_{p}$ norms are equivalent.
Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a reasonable privacy-utility trade-off under pure $\varepsilon$-DP for private fine-tuning of LLMs. While this work primarily explores these two mechanisms, the DP-ZO framework is flexible enough to be extended to other differential privacy mechanisms, broadening its applicability.
Privacy Analysis.
As we consider multiple accounting methods with multiple previously proposed mechanisms, we give the overview of the analysis below and defer the full privacy analysis to Appendix A.
Theorem 3.1.
Algorithm 1 is $(\varepsilon,\delta)$-DP.
Proof Overview. We analyze the privacy of models trained iteratively with DP-ZO by composing the per-iteration privacy loss (Kairouz et al., 2015). At each step, line 11 upper bounds the $\ell_{1}$ (and therefore $\ell_{p}\forall p\geq 1$) sensitivity of the difference in losses $l_{i}$ by $C$. Line 12 adds noise based on DP mechanisms such that each step in Algorithm 1 satisfies $(\varepsilon,\delta)$-DP with some privacy parameters. Therefore Algorithm 1 is $(\varepsilon,\delta)$-DP with some privacy parameters that are calculated via some mechanism-dependent composition theorem. The specific $\varepsilon,\delta$ depends on the choice of mechanism. Most of the privacy analyses in Appendix A are based on the numerical composition of private random variable for the corresponding privacy curve (Gopi et al., 2021) except for the pure $\varepsilon$-DP analysis. The analysis for pure $\varepsilon$-DP by the Laplace mechanism is based on the basic composition (Dwork and Roth, 2014). Given a number of iterations $T$ and the use of the Laplace mechanism $Laplace(0,\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
Remark. We can get a tighter privacy composition of $\varepsilon$ by relaxing the Laplace mechanism’s $\varepsilon$-DP guarantee to $(\varepsilon,\delta)$-DP guarantee. We can compute the composition of privacy loss for Laplace mechanism by the private random variable for privacy curve of a $\varepsilon$-DP algorithm in Gopi et al. (2021). We also provide a more fine grained privacy analysis by using the dominating pair for the subsampled Laplace mechanism in the accounting method of Gopi et al. (2021) (See Appendix A).
Table 1: Main results with 1000 training samples for each dataset. OPT-13B model with LoRA fine-tuning. DP-ZO (G) is DP-ZO instantiated with the Gaussian mechanism. $\delta=10^{-5}$. The $\varepsilon=\infty$ by ZO is $84.73$ for SQuAD, $30.78$ for DROP, and $92.09$ for SST2. The $\varepsilon=0$ baseline, i.e., directly doing model evaluation without training, is $46.23$ for SQuAD, $14.64$ for DROP, and $58.83$ for SST2.
Task SQuAD DROP SST2
Task type generation (metric: F1) classification (metric: accuracy)
Method $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$ $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$ $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$
DP-ZO(G) $79.22_{0.64}$ $81.35_{0.38}$ $82.85_{0.31}$ $28.27_{0.40}$ $29.45_{0.40}$ $30.34_{0.54}$ $85.07_{3.31}$ $89.68_{0.82}$ $92.00_{0.94}$
Table 2: DP-ZO (Gaussian) and DP-SGD with full parameter and LoRA fine-tuning on SQuAD with 1000 training samples across different model sizes. $(1,10^{-5})$-DP. ‘-’ means the approach did not scale with straightforward implementation;  Section 5 details the additional engineering required to scale DP-SGD to larger models. ‘- -’ for DP-ZO means the results are omitted due to limited computational resources.
Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B OPT-66B
DP-ZO-LoRA (Gaussian) $74.41_{0.81}$ $79.30_{0.93}$ $79.53_{0.65}$ $81.35_{0.38}$ $82.03_{0.90}$ $83.87_{0.99}$
DP-SGD-LoRA $75.39_{0.33}$ $79.42_{0.57}$ $79.53_{0.52}$ $82.14_{0.18}$
DP-ZO-Full (Gaussian) $72.84_{1.03}$ $77.25_{0.27}$ $79.06_{0.67}$ $82.16_{0.41}$
DP-SGD-Full $75.50_{0.89}$ $79.81_{0.64}$
4 EVALUATION
We first overview our experimental setup in Section 4.1 and then evaluate the performance of DP-ZO in Section 4.2. We find that DP-ZO provides a competitive privacy-utility tradeoff for conservative privacy budgets across multiple datasets and model scales. One noteworthy result is that DP-ZO exhibits just $1.86\%$ accuracy degradation due to privacy at $\varepsilon=1$ when fine-tuning OPT-66B on 1000 training samples from SQuAD. We also compare DP-ZO to DP-SGD in Section 4.2. In Section 4.3 we first characterize DP-ZO under different few-shot settings. We then show that DP-ZO is robust to different mechanisms for noise addition. Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP. We also discuss the choice of batch size for DP-ZO.
4.1 Experimental Setup
When we report results, we report the metric of interest (F1 score or accuracy) and standard deviation averaged across 5 independent runs with different random seeds. We detail the full hyperparameter searches in Appendix D.
Datasets.
We consider three different benchmark NLP tasks: SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation, and SST2 (Socher et al., 2013) for text classification. Although all three datasets have very different dataset sizes, we consider the few-shot setting for all three datasets where we sample $1000$ points for each dataset. Fine-tuning LLMs with $n$ samples on the order of $n=1000$ is a standard setting widely studied in the NLP community (Gao et al., 2021; Malladi et al., 2023) because we are generally interested in the few-shot abilities of LLMs (Brown et al., 2020). This represents a departure from prior works that privately finetune LLMs; Yu et al. (2022); Li et al. (2022b); Yu et al. (2021b) use the entire training dataset of SST2 that has about 65,000 examples. It is well known that the privacy-utility tradeoff improves greatly with more data (Tramèr and Boneh, 2021). It is straightforward to see that our setting with datasets of the size $n=1000$ with $\delta=10^{-5}$ is simultaneously more challenging and more aligned with real-world usecases than previous works in DP finetuning of LLMs. Despite the increased difficulty of our few-shot setting as compared to prior work, our results validate that DP-ZO realizes a strong privacy-utility tradeoff.
Models.
We use models from the OPT (Zhang et al., 2022) family of decoder-only transformers. We present our main results (Table 1) using a pretrained OPT-13B model that is finetuned with LoRA (Hu et al., 2022); that is, we update $<1\%$ of the total parameters. We include a range of ablation studies, including varying the model size and amount of parameters to be updated after we present main results.
Privacy Budgets.
We consider various privacy levels with $\varepsilon=[0.5,1,4]$ and fix $\delta=10^{-5}$ for $(\varepsilon,\delta)$-DP. We include the zero-shot $\varepsilon=0$ baseline does not incur any privacy loss because we evaluate the pretrained model directly without finetuning on private data. We also include the non-private $\varepsilon=\infty$ baseline that is trained without any DP guarantee. That is, we iterate over the shuffled dataset instead of doing Poisson sampling (replacing line 4), do not clip the step size (skipping line 11) and set $\sigma=0$ (in line 12). We make these modifications between Poisson sampling and clipping are known to degrade performance, and we want to compare to the strongest possible nonprivate baseline.
4.2 Main Results
DP-ZO Provides a Strong Privacy-utility Tradeoff for Conservative Privacy Budgets.
As shown in Table 1, across all three tasks and all $\varepsilon$s, DP-ZO significantly improves upon the $\varepsilon=0$ baseline, and only slightly degrades the performance compared to the non-private baseline. For SQuAD, even at $\varepsilon=0.5$, DP-ZO can still achieve $78.10\%$, that significantly outperforms $\varepsilon=0$ baseline ($46.23\%$). The gap between $\varepsilon=0.5$ and $\varepsilon=\infty$ is about $5.5\%$. By increasing $\varepsilon$ from $0.5$ to $4$, this gap can be further reduced to $2\%$. For DROP and SST2, DP-ZO (Gaussian) achieves comparable performance as the non-private baseline at $\varepsilon=4$.
DP-ZO Scales to Large Models.
In Table 2 we show that DP-ZO continues improving as the model size increases from 1.3B to 66B. Due to space constraints, we provide the non-private ($\varepsilon=\infty$) performance of all models and methods in Appendix E. Table 2 shows an promising insight: as the model size and nonprivate performance increase, the gap in performance between private and nonprivate models shrinks. Specifically, the gap for OPT-125M is $4.96\%$ ($48.7\%$ at $\varepsilon=\infty$ reduced to $43.74\%$ under $\varepsilon=1$; omitted from Table 2 for poor performance). But this gap shrinks to just $1.86\%$ for OPT-66B, where the private performance at $\varepsilon=1$ is $83.87\%$ compared to $85.73\%$ non-privately. Our findings therefore suggest that DP-ZO scales to large models not only because it is compatible with existing pipeline without much additional engineering effort but also because the utility drop due to privacy is smaller as the model size increases.
Comparison with DP-SGD.
We compare DP-ZO to differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016) which has been applied to fine-tune LLMs with full parameter fine-tuning (Li et al., 2022b) and with LoRA (Yu et al., 2022; He et al., 2023). Recall that DP-ZO is compatible out-of-the-box with mixed precision training and GPU parallelism, enabling us to fine-tune OPT-66B. As we show in  Section 5, it is significantly more challenging to integrate DP-SGD with these techniques, and furthermore, DP-SGD requires more memory than DP-ZO to store activations and compute per-sample gradients. As a direct result, DP-SGD cannot directly scale past 2.7B with full fine-tuning or 13B with LoRA without additional implementations desired for multiple GPUs, while DP-ZO can scale seamlessly to larger models. In Table 2 we present comparisons between DP-ZO and DP-SGD with full parameter finetuning and LoRA. The best performance by DP-ZO is  83.87% by OPT-66B finetuned with LoRA. This is $\approx 2\%$ better than the best performance of DP-SGD in Table 2 that is $82.14\%$ by OPT-13B with LoRA. Note that when compared to the same model size, the performance of DP-ZO is $1\%$ worse than DP-SGD with LoRA and $2-3\%$ worse than DP-SGD with full parameter funetuning. How to close the gap between DP-ZO and DP-SGD for the same model size is an interesting direction and we leave it as future work.
4.3 Analysis
We vary the amount of training data we sample, the choice of DP mechanism in DP-ZO, and the batch size.
Table 3: Ablation of DP-ZO (Gaussian) for different $n$ training samples on SQuAD dataset. ($1,10^{-5}$)-DP. OPT-13B with LoRA finetuning.
$n$-shot $n=250$ $n=500$ $n=1000$ $n=5000$
$\varepsilon=1$ $73.86_{0.76}$ $77.43_{0.51}$ $81.35_{0.38}$ $82.35_{0.48}$
$\varepsilon=\infty$ $84.03$ $84.66$ $84.73$ $84.75$
Characterizing Differentially Private Few-Shot Learning.
Although it is known that private learning requires more data than non-private learning (Bassily et al., 2014), prior work has not characterized this improvement for fine-tuning language models. In Fig. 1 and Table 3 we vary the number of training samples $n$ around the $n=1000$ setting in the main results while keeping $\delta=10^{-5}$ fixed for all choices of $n$ such that the largest $n=5000$ still satisfies $\delta\ll 1/n$.
Figure 1: (Smoothed) training loss. $n=5000$ has better convergence rate compared to $n=250$.
Table 3 shows that DP-ZO can achieve nontrivial performance in few-shot settings under conservative privacy guarantees. Furthermore, we find that while increasing the amount of training data by 10×10$\times$10 × barely increases non-private performance, it increases private performance by $\approx 5\%$. While non-private few-shot learning can succeed by just memorizing the training data, Fig. 1 indicates that the convergence rate for different shots for private few-shot learning is different. With the proliferation of pretrained models, we anticipate that privately fine-tuning downstream tasks in the few-shot setting will be more aligned with real-world use cases.
Table 4: Pure $\varepsilon$-DP by DP-ZO (Laplace), SQuAD with 1000 training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ performance with ZO is $84.73\%$ and the $\varepsilon=0$ baseline is $46.23\%$.
$\varepsilon$ $\varepsilon=4$ $\varepsilon=10$ $\varepsilon=15$
DP-ZO (Laplace) $64.50_{1.04}$ $73.69_{1.25}$ $76.99_{0.76}$
DP-ZO with Pure $\varepsilon$-DP.
To the best of our knowledge, DP-ZO (Laplace) is the first method that achieves a non-trivial privacy-utility tradeoff under pure $\varepsilon$-DP. In Table 4 we find that DP-ZO (Laplace) can significantly improve upon $\varepsilon=0$, albeit for a large privacy budget. Given a budget $\varepsilon=10$, which some prior work has considered reasonable (Ponomareva et al., 2023), DP-ZO (Laplace) can obtain $73.69\%$ on SQuAD. As we showed in Table 3, the few-shot setting poses a unique challenge for obtaining strong performance under conservative privacy budgets. Acquiring more data enhances privacy amplification and reduces the amount of noise we need to add to achieve a target $\varepsilon$-DP guarantee (Figure 2). In particular, in Table 5 we find that increasing the number of training examples from $1000$ to $5000$ improves performance at $\varepsilon=4$ from $64.50\%$ to $78.65\%$, although the improvement of non-private performance at $\varepsilon=\infty$ by increasing training samples from $1000$ to $5000$ is insignificant. DP-ZO (Laplace) presents a new direction for obtaining privacy guarantees without failure probabilities.
Table 5: Pure $\varepsilon$-DP by DP-ZO (Laplace) at $\varepsilon=4$, SQuAD with different training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ by ZO is $84.73\%$ and $84.75\%$ for 1000 and 5000 samples respectively. The $\varepsilon=0$ baseline is $46.2\%$.
$n$-shot $n=1000$ $n=5000$
DP-ZO (Laplace) $64.50_{1.04}$ $78.65_{1.37}$
Figure 2: The added noise scale $\sigma$ for different number of samples to satisfy $(4,0)$-DP decreases with data size.
Table 6: DP-ZO with different DP mechanism. SQuAD with 1000 training samples. $\delta=10^{-5}$. G is for Gaussian and L is for Laplace.
$\varepsilon$ $\varepsilon=0.35$ $\varepsilon=0.5$ $\varepsilon=1$ $\varepsilon=4$
DP-ZO (G) $76.90_{0.23}$ $79.22_{0.64}$ $81.35_{0.38}$ $82.85_{0.31}$
DP-ZO (L) $73.69_{1.25}$ $76.76_{0.74}$ $80.10_{0.85}$ $82.70_{0.39}$
Different Noise Mechanisms for $(\varepsilon,\delta)$-DP.
We now relax the privacy guarantee provided by the Laplace mechanism to approximate $(\varepsilon,\delta)$-DP. For example, the privacy parameter setting given by basic composition as $(10,0)$-DP in Table 4 can also be tracked via composition of private random variable for privacy curve of a $\varepsilon$-DP algorithm and costs $(0.35,10^{-5})$-DP. In Table 6, we compare DP-ZO instantiated with the Laplace and Gaussian mechanisms. DP-ZO (Gaussian) outperforms DP-ZO (Laplace) for strict privacy budgets such as $\varepsilon=0.35$ because it enjoys tighter accounting (Gopi et al., 2021) and lower variance (Dwork and Roth, 2014). These advantages are less significant for larger privacy budgets; for $\varepsilon=4$, both DP-ZO (Gaussian) and DP-ZO (Laplace) exceed $82\%$.
Ablation Study of Batch Size for DP-ZO.
The tradeoff between decreasing the batch size to benefit from privacy amplification by subsampling (Balle et al., 2018) and increasing batch size to minimize the per-example added noise (De et al., 2022; Ponomareva et al., 2023) has trended towards training with large batches (Li et al., 2022b) or even doing full-batch gradient descent (Panda et al., 2022). However, these analyses are specific to SGD, and the optimal batch size for DP-ZO is of course an open question.
Figure 3: (Smoothed) training loss. Larger batch sizes have a slightly better convergence rate and the convergence rates for batch size of 16, 32, 64 are similar.
In Fig. 3 and Table 7 we ablate the batch size $B$ in DP-ZO (Gaussian) around the batch size $B=16$ we use for our main results. Somewhat surprisingly, we find that increasing the batch size does not strictly improve performance for $(1,10^{-5})$-DP, even though it has been shown to improve performance in the non-private setting (Liu et al., 2018).
Table 7: Different batch size $B$ on SQuAD 1000 training samples. OPT-13B model. $\varepsilon=1$. 10,000 steps.
$B$ 4 8 16 32 64
$\sigma$ $1.68$ $3.11$ $6.08$ $12.07$ $24.09$
F1 $79.94$ $80.95$ $81.35$ $81.63$ $81.47$
Std. $0.29$ $0.63$ $0.38$ $0.62$ $0.81$
One possible reasoning for this is the nature of zeroth-order optimization. Under DP-SGD, increasing the batch size decreases the variance of the gradient estimator and therefore decreases the number of iterations required for convergence (Bassily et al., 2020), which in turn reduces the amount of noise that must be added at each iteration of DP-SGD to ensure privacy (Kairouz et al., 2015) and counterbalances the loss of amplification. Note that this does not reduce the actual amount of computation done, because DP-SGD typically requires gradient accumulation to train large models with large batch sizes (De et al., 2022). By contrast, because SPSA inherently has higher variance than SGD (Malladi et al., 2023), the variance reduction by large batch size in DP-ZO may not be as significant as on DP-SGD. An interesting direction for future work is designing a batch size schedule around the intricacies of SPSA. In Appendix D we did an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
5 DISCUSSION
In Section 4.2 we showed that DP-ZO obtains superior results to DP-SGD by scaling to larger models. Now we examine the amount of engineering effort necessary to scale DP-SGD to larger models, a topic on which many papers have been written (Bu et al., 2023a, b; Yousefpour et al., 2021; Li et al., 2022b; He et al., 2023). We find that DP-ZO seamlessly scales to larger models and believe its simplicity presents a compelling alternative to DP-SGD for practitioners.
DP-SGD.
Differentially Private Stochastic Gradient Descent (DP-SGD) (Song et al., 2013; Abadi et al., 2016) is the standard privacy-preserving training algorithm for training neural networks on private data, with an update rule given by $w^{(t+1)}=w^{(t)}-\frac{\eta_{t}}{|B|}\left(\sum_{i\in B}\frac{1}{\text{c}}\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))+\sigma\xi\right)$ where the changes to SGD are the per-sample gradient clipping $\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))=\frac{c\times\nabla\ell(x_{i},w^{(t)})}{\max(c,||\nabla\ell(x_{i},w^{(t)})||_{2})}$ and addition of noise sampled from a $d$-dimensional Gaussian distribution $\xi\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. DP-SGD is the marquee algorithm for privacy-preserving machine learning, but it requires implementing per-example gradient clipping. This creates a slew of challenges for deploying DP-SGD.
Computational and Memory Challenges in DP-SGD. DP-SGD requires the computation of per-example gradients, which can be naively implemented by storing each gradient in the batch separately. This approach inflates the memory overhead by a factor of $B$, where $B$ is the batch size. Tensorflow Privacy avoids this issue by clipping microbatches rather minibatches, which does not slow down training but increases the noise added and therefore hurts utility. Jax can automatically vectorize the per-sample gradient computation, but training is still slowed down. Recently, specialized libraries have been developed that instead analytically compute the norm of the gradients for different layers (Li et al., 2022b; Bu et al., 2023c). This requires actually implementing the computation, which is challenging for new layers. If a practitioner wants to train LLaMA2 and efficiently compute the per-example gradient norms, they would first have to derive the analytical formula for the norm of GroupedQueryAttention (Touvron et al., 2023), which can represent a nontrivial amount of engineering. This complexity is further compounded when considering models with parameter sharing, such as the OPT networks we use in our experiments, as this is generally not compatible with analytical norm clipping methods (lxuechen, 2022). Although network architectures exist that do not require parameter sharing, the best pretrained models (Zhang et al., 2022; Touvron et al., 2023) use parameter sharing because it makes the model use the same representation for predicting tokens at the next step as well as for decoding.
5.1 DP-ZO Scales Seamlessly
In order to train models as large as OPT-66B, whose parameters cannot be loaded into memory on a single A100 80GB GPU, we need to implement some form of parallelism across GPUs. We now discuss how easy this is for DP-ZO (in the simplest form, just running DP-ZO on a machine with 2 GPUs will prompt HuggingFace to implement naive model parallelism) and how challenging it can be for approaches that require per-example gradient clipping such as DP-SGD.
Data Parallelism in DP-ZO. To synchronize model state between GPUs in data-parallel-DP-ZO, we just transfer the random seed and its corresponding fp16 scalar step size; this is just a few bytes. However, first-order approaches such as DP-SGD require the transfer of gradients across devices to update all the models, necessitating expensive allgather and reduce operations. This communication overhead is $1.5d$ in PyTorch FSDP, where $d$ is the size of the model.
DP-ZO Does Not Store Gradients. DP-ZO does not store activations or gradients in the forward pass, thus conserving memory. DP-SGD needs to store the activations at each GPU under pipeline parallelism in order to clip the per-example gradient (He et al., 2023), which will fill up the GPU memory and limit new microbatches from being processed.
DP-ZO Easily Integrates With Mixed Precision Training. All of our DP-ZO experiments use mixed precision; we load the model and compute the loss in half precision. Mixed precision training is much faster than full precision and also has a smaller memory footprint; these factors contribute to DP-ZO’s ability to train much larger models. By contrast, implementing DP-SGD with half precision is challenging. Yu et al. (2022) do not modify their implementation of DP-SGD to support half precision and report worse performance as a result. Li et al. (2022b) detail an algorithm that interleaves loss scaling with gradient clipping to avoid underflow while maintaining utility. This implementation is nontrivial and as of today is not available in Opacus (ffuuugor, 2022).
DP-ZO is Storage and Communication-Efficient Even After Training Has Completed. DP-ZO offers significant advantages in terms of storage and communication efficiency, especially beneficial for bandwidth-constrained environments like edge devices. Unlike traditional methods where the difference in model parameters $\theta_{0}-\theta_{f}$ is shared—which could amount to multiple gigabytes for large models—DP-ZO allows for the storage and transmission of a sequence of updates. This sequence is represented as an array of tuples $[(\textsc{seed}_{0},0.54),\cdots,(\textsc{seed}_{f},-0.14)]$, where each tuple contains a seed and a step size, taking up only 4 bytes. Even for $1\times 10^{4}$ fine-tuning iterations, this array would require less than 1MB of storage, representing a substantial reduction in both storage and communication overhead. We can apply these weight differences to a model by simply iterating over the array, sampling from the PRNG using the given seed, scaling that random vector, and applying it to the current model parameters. This procedure is highly efficient, as it involves only sequential memory accesses and scalar floating-point operations.
6 RELATED WORK
In this section we give an overview of the broader body of work privacy preserving large language models and private zeroth-order optimization method.
Privacy Preserving Large Language Models.
Recent studies have leveraged DP-SGD to fine-tune large language models. Li et al. (2022b) provide methods for fine-tuning large language models with DP-SGD by ghost clipping to mitigate the memory burden of per-sample gradient clipping. Yu et al. (2022) report compelling results by only updating a sparse subset of the LLMs with parameter efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2022). He et al. (2023) leverage group-wise clipping with adaptive clipping threshold and privately fine-tune the 175 billion-parameter GPT-3. Duan et al. (2023); Li et al. (2022b) also consider private prompt tuning by adding noise to the soft prompt (Li and Liang, 2021; Lester et al., 2021). Du et al. (2023) add non-i.i.d. noise from a matrix Gaussian distribution to directly perturb embedding in the forward pass of language models. With the emergence in-context learning of large language models (Brown et al., 2020), recent works (Duan et al., 2023; Wu et al., 2023; Tang et al., 2023b) study private in-context learning of large language models without fine-tuning.
Private Zeroth-order Optimization.
Most recently, a concurrent work (Zhang et al., 2023) also considers the same DP-SPSA algorithm for zeroth-order optimization. Our method and Zhang et al. (2023) are functionally the same up to constants, and our work focuses on an empirical evaluation of the method, whereas Zhang et al. (2023) extends the convergence analysis of Malladi et al. (2023) to DP as shown in Appendix B of Zhang et al. (2023). There is a slight difference for the generation of random perturbation of Zhang et al. (2023) and our Algorithm 1. Zhang et al. (2023) uses the random unit vector for the perturbation and the convergence analysis is based on such set-up, whereas our perturbation is a normally distributed vector. Note that Algorithm 1 and 2 in Zhang et al. (2023) also scales the unit vector by the square root of the model dimension, so the two approaches are functionally the same. We reimplemented our perturbation method based on the algorithms in Zhang et al. (2023), and we obtain the same performance of $81.3\%$ F1 with $(1,10^{-5})$-DP for OPT-13B by DP-ZO by LoRA finetuning on SQuAD.
Gratton et al. (2022) analyze the intrinsic privacy of the zeroth-order optimization for DP-ADMM (Huang et al., 2020) in distributed learning.  Gratton et al. (2022) state that if the output of the zeroth-order method itself follows Gaussian distribution, the noise can be analyzed as the Gaussian mechanism and provide intrinsic privacy. However, this is merely stated as an assumption for lemma 1. To the best of our knowledge there is no work that proves that the zeroth-order gradient estimator can actually be analyzed as the sum of an unbiased gradient estimator and some Gaussian error term.
7 CONCLUSION
DP-SGD has been the de-facto private training method of the last decade. In this work we propose DP-ZO, a novel method for private fine-tuning that privatizes the zeroth-order update by adding noise to the difference in loss between two perturbations. DP-ZO’s unique univariate privatization unlocks training larger models with better parallelism than DP-SGD. In particular, DP-ZO can train OPT-66B at $(1,10^{-5})$-DP in the challenging few-shot setting with only $1.86\%$ performance degradation due to privacy, beating DP-SGD by $1.71\%$. We also achieve the first non-trivial pure $\varepsilon$-DP result by instantiating DP-ZO with the Laplace mechanism. We anticipate that future work can further study these design choices, integrate more DP mechanisms into DP-ZO, and apply it to the vision domain.
Acknowledgement
This work was supported in part by the National Science Foundation under grant CNS-2131938, the ARL’s Army Artificial Intelligence Innovation Institute (A2I2), Schmidt DataX award, and Princeton E-ffiliates Award.
Appendix A Privacy Analysis
Proposition A.1 (Basic Composition theorem (Dwork and Roth, 2014)).
If $M_{1}$ is $(\varepsilon_{1},\delta_{1})$-DP and $M_{2}$ is $(\varepsilon_{2},\delta_{2})$, then the adaptive composition of $M_{1}$ and $M_{2}$ is $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP.
Proposition A.2 (Privacy Amplification via Subsampling (Balle et al., 2018)).
If M is $(\varepsilon,\delta)$-DP, then the subsampled mechanism with sampling rate $p$ obeys $(\varepsilon^{\prime},\delta^{\prime})$-DP with privacy parameters $\varepsilon^{\prime}=\log(1+p(e^{\varepsilon}-1))$ and $\delta^{\prime}=p\delta$.
DP-ZO can be instantiated with different noise mechanisms. In this subsection we provide privacy analysis for the Gaussian mechanism and Laplace mechanism.
Gaussian Mechanism
As outlined in Line 11, the $\ell_{2}$ sensitivity of Alg. 1 is $\mathrm{C}$ and we are adding $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ noise to the estimated loss. We analyze the composition of subsampled Gaussians with the privacy loss variable accounting approach of Gopi et al. (2021).
Laplace Mechanism
We first analyze the privacy cost for one step by the Laplace mechanism. At each step, we sample a new batch of data with the sample rate of $p=B/|\mathcal{D}|$. As outlined in 11, the $\ell_{1}$ sensitivity of Algorithm 1 is $\mathrm{C}$. By Proposition 2.3, the privacy cost at one step would cost $(1/\sigma,0)$-DP on this batch. By Proposition A.2, the privacy cost at one step would cost $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP on the full dataset $\mathcal{D}$. By  Proposition A.2, the privacy cost of Algorithm 1 instantiated with Laplace mechanism satisfies $(T\cdot\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP. We provide the privacy parameters we used for pure $\varepsilon$-DP by the Laplace mechanism in Table 9.
We can also get tighter composition of $\varepsilon$ with relaxation to $\delta>0$. Specifically, we can instantiate the PRV of random response with $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$ because the dominating pair for random response is a dominating pair for the pure DP mechanism. Then, we can use the numerical composition of Pure DP PRV by Gopi et al. (2021). We use this privacy accounting method for the approximate $(\varepsilon,\delta)$-DP results for Laplace mechanisms in the main paper.
In addition, we could also compute the privacy cost of composition for the Laplace mechanism by Monte Carlo based DP accountant (Wang et al., 2023). Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Let $b$ be the scale of Laplace noise, $p$ the sub-sampling rate, and assume the sensitivity is $1$, and assume we are doing composition for $T$ iterations, each iteration with sampling rate $p$. By Zhu et al. (2022) we know that the pair of distribution $(P,Q)$ dominating pair for a single dimensional Laplace mechanism, where $P$ and $Q$ are distributed according to the following pdfs,
$f_{P}=\frac{1}{2b}\exp(-|x|/b)~{}~{}~{}\text{and}~{}~{}~{}f_{Q}=\frac{1}{2b}\exp(-|x-1|/b).$
Therefore, ($P,(1-p)\cdot P+p\cdot Q$) is the dominating pair for the sub-sampled Laplace. We plug this into the standard Monte-Carlo accountant of Wang et al. (2023) (without importance sampling, see Algorithm 2 and Theorem 10 in Wang et al. (2023)) while using $10^{10}$ samples to calculate the $\delta$ at a given value of $\varepsilon$. Also, using the analytical accountant explained above, we always make sure that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $10^{-8}$ (We use the fact that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $\mathbb{E}[PRV^{2}]$ and the fact the $PRV$ is always bounded for Laplace mechanism.). This ensures that the error of our estimation of $\delta$ is at most $10^{-8}$ with probability at least $1-10^{-5}$. Putting all together, for all reported values of $\varepsilon$, our $\delta$ is bounded by $10^{-5}$, with probability at least $0.99999$. This privacy analysis is tighter with $\varepsilon$ is high compared to the former analysis which use the pure DP PRV accountant. This is consistent with the intuition. As we increase the distance between the Laplace dominating pairs, the probability of sampling points from the area between the centers increases. And that is where the Laplace Mechanism is different from the Randomized Response. We present the accounting results for the Laplace method to achieve $(\varepsilon,\delta)$-DP by these two accounting methods in Table 9. Table 9 shows that the Monte Carlo based DP accountant can give tighter analysis for the Laplace mechanism for $(\varepsilon,\delta)$-DP than the pure $\varepsilon$-DP PRV method.
Table 8: $\varepsilon$-DP by Laplace. BSZ=16, Steps=10000.
Table 9: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=10000.
$\sigma$ $|D|$ $\varepsilon$
40.5 1000 4
16.5 1000 10
11.2 1000 15
8.49 5000 4
$\sigma$ $\varepsilon$ (by Monte-Carlo) $\varepsilon$ (by pure-DP PRV)
16.5 0.33 0.35
11.76 0.48 0.5
6.5 0.9 1
2.19 $2.9$ 4
Table 9: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=10000.
Appendix B Implementation Details
We follow Malladi et al. (2023) and provide the memory-efficient version of DP-ZO in Algorithm 2. Algorithm 2 enjoys the benefit that it does not incur additional GPU memory cost compared to inference.
Algorithm 2 Differentially Private-ZO (GPU memory efficient version. Adapted from Malladi et al. (2023))
1:Model parameters $\theta$, dataset $\mathcal{D}$, learning rate $\alpha$, perturbation scale $\phi$, random seed $s$, weight decay $\lambda$, noise scale $\sigma$, noising mechanism $\mathcal{Z}$, clipping threshold ${C}$, expected batch size $B$ and sampling rate $p=B/|\mathcal{D}|$. Lines with * are DP modifications.
2:procedure DP-ZO($\theta$, $\mathcal{D}$, $\varepsilon$, $\sigma$, $T$, $s$, $\phi$, $C$, $\alpha$)
3:     for $t\in 1,\ldots T$ do
4:         Poisson samples $\mathcal{B}$ from dataset D with sampling rate $p$
5:         $\theta\leftarrow$ PerturbParameters($\theta,\phi,s$)
6:         Compute per-sample loss $\mathcal{L}_{1}(\theta,\mathcal{B})$
7:         $\theta\leftarrow$ PerturbParameters($\theta,-2\phi,s$)
8:         Compute per-sample loss $\mathcal{L}_{2}(\theta,\mathcal{B})$
9:         $\theta\leftarrow$ PerturbParameters($\theta,\phi,s$)
10:         Compute difference in loss $\mathcal{L}=\mathcal{L}_{1}-\mathcal{L}_{2}$
11:         Clamp $\mathcal{L}$ between $-{C}$ and ${C}$
12:         $g=\frac{\sum_{i\in\mathcal{B}}L+\mathcal{Z}({C},\sigma)}{B*2\phi}$
13:         Reset random number generator with seed $s$
14:         for $\theta_{i}\in\theta$ do
15:              $z\sim\mathcal{N}(0,1)$
16:              $\theta_{i}\leftarrow\theta_{i}-\alpha*g*z$
17:procedure PerturbParameters($\theta,\phi,s$)
18:     Reset random number generator with seed $s$
19:     for $\theta_{i}\in\theta$ do
20:         $z\sim\mathcal{N}(0,1)$
21:         $\theta_{i}\leftarrow\theta_{i}+\phi z$
Appendix C Design Choices
Algorithm 1 outlines our DP-ZO that estimates the gradients via privatized loss value without backpropagation. In this subsection, we provide several design choices for Algorithm 1.
Definition 2 (n-SPSA Gradient Estimator) The n-SPSA gradient estimate averages $\nabla L_{b}(\theta;B)$ over $n$ randomly sampled $z$. We can write this in vector notation, dropping the normalizing constants for succinctness.
$g_{i}$ $=L(\theta+\varepsilon z_{i};B)-L(\theta-\varepsilon z_{i};B)\text{(projected gradient for each }i)$
$\mathbf{Z}$ $=[z_{1},z_{2},...,z_{n}]\text{(matrix whose columns are the }z\text{ vectors)}$)
[…]
Then the n-SPSA gradient estimate can be written as:
$\begin{split}\nabla L_{n}(\theta;B)=\mathbf{g}\cdot Z\quad(2)\end{split}$
How Many Gradients to be Estimated in a Model Update.
Algorithm 1 estimates the gradients once. As outlined above, SPSA can be extended to n-SPSA gradient estimator and n-SPSA can improve the performance in the non-private setting (Malladi et al., 2023). Here we discuss our design choice of why we choose $n=1$ in Algorithm 1.
  • Estimate the average. Previous work (Malladi et al., 2023) shows that averaged estimation helps the non-private setting. In a private setting, we have to privatize the gradient estimation. Here we discuss our initial design of the privatized n-SPSA gradient estimation. For the sampled batch, assuming we are adding the Gaussian noise $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ for 1-SPSA. Then for n-SPSA, to ensure we have the same privacy cost as 1-SPSA, we need to add $\mathcal{N}(0,n\cdot\mathrm{C}^{2}\sigma^{2})$ to each gradient estimation and finally average the $n$ gradients. Our privacy analysis follows the $n$-fold composition of Gaussian mechanism (Corollary 3.3 in Gaussian differential privacy (Dong et al., 2019)). Our initial experiment result shows that our current analysis for n-SPSA noise addition does not make n-SPSA improve in the private setting compared to 1-SPSA. We leave the improvement in tighter analysis for private n-SPSA as future work and use 1-SPSA to conduct experiments.
The Type of Noise for DP.
As discussed in Section 3Algorithm 1 can be incorporated in different noise mechanisms. We focus on the Gaussian noise mechanism and the Laplace mechanism in this work. The Gaussian noise mechanism has been widely studied in previous literature both for privacy analysis and empirical performance in DP-SGD (Abadi et al., 2016; Mironov, 2017; Dong et al., 2019). The Laplace mechanism, though less studied for privacy-preserving machine learning, can provide pure DP while the Gaussian mechanism can only provide approximate DP. We have provided the privacy analysis in Section A.
Appendix D Hyperparameter Search
We have presented the hyperparameter search for batch size in Table 7. In this section, we first provide several initial results for hyperparameter search on clipping threshold and finally present the hyperparameter tables. We also provide an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
Different Clipping Threshold.
Li et al. (2022b); De et al. (2022) recommend small clipping $C$ threshold for DP-SGD training. For example, Li et al. (2022b) use $C=0.1$ for training language models. We therefore study the effect of different clipping threshold and present the results in Table 10. We find that while $C=1$ performs significantly worse, setting $C$ as $0.1$, $0.05$, $0.01$ are within the 2% performance gap. We therefore choose $C=0.05$.
Table 10: Different clipping C. $\sigma=15.9$. batch size=16, 10,000 steps. $\varepsilon=0.35$.
Clip=1 Clip=0.1 Clip=0.05 Clip=0.01
F1 $66.04$ $74.26$ $76.81$ $75.39$
Hyperparameter for DP-ZO (Gaussian) in Main Results.
We present the hyperparameter for DP-ZO (Gaussian) in Table 11.
Table 11: Hyperparameter search for DP-ZO in main results Table 1.
$|\mathcal{D}|$ 1000
Steps $T$ $10000$
Clipping $C$ $0.05$
Batch size $16$
$\sigma$ $11.47$ for $\varepsilon=0.5$, $6.08$ for $\varepsilon=1$, $1.88$ for $\varepsilon=4$
learning rate [2e-5, 5e-5, 1e-4, 2e-4, 5e-4] for LoRA finetuning.
[2e-7, 5e-7, 1e-6, 2e-6, 5e-6] for Full fine-tuning.
LoRA rank 8
Hyperparameter for DP-SGD.
We present the hyperparameter search for DP-SGD in Table 12.
Table 12: Hyperparameter search for DP-SGD in Table 2.
$|\mathcal{D}|$ 1000
Steps $T$ $200$
Clipping $C$ $0.1$
Batch size $64$
$\sigma$ $6.60$ for $\varepsilon=0.5$, $3.59$ for $\varepsilon=1$, $1.28$ for $\varepsilon=4$
learning rate [1e-4, 2e-4, 5e-4, 1e-3, 2e-3] for LoRA fine-tuning.
[1e-5, 2e-5, 5e-5, 1e-4, 2e-4, 5e-4] for Full fine-tuning.
LoRA rank 8
Hyperparamter for DP-ZO (Laplace).
The hyperparameter search for DP-ZO (Laplace) is similar to DP-ZO (Gaussian).
Ablation on Batch Size and Steps.
In Fig. 5 and Fig. 5, we did an initial study to systematically evaluate the interplay between batch size and training iterations by varying batch size in [4,8,16,32,64] and steps in [10000, 2000]. Similar to main results, we run 5 independent runs for each setting and compute the average of 5 runs.
Figure 4: Ablationg ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples.
Figure 4: Ablationg ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples.
Figure 5: Ablationg DP-ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples. $(1,10^{-5})$-DP.
Appendix E Ablation on Model Size
Section 4.2 shows that DP-ZO scales to larger models and provides the results of DP-ZO for model size varying from 1.3B to 66B parameters in Table 2. Here we provide the full results of DP-ZO finetuned with LoRA at $\varepsilon=1$, with model size ranging from 125M to 66B. We also include the $\varepsilon=0$ and $\varepsilon=\infty$ baseline as a reference in Table 13.
Table 13 shows the full trend of DP-ZO with model size scaling from 125M to 66B, that is DP-ZO scales to larger models.
For OPT-125m, the gap between private and non-private baseline is $4.96$. For OPT-66B, the non-private baseline is 85.73 and the gap between the private and non-private results is $1.86$.
Table 13: Ablation of DP-ZO across different model sizes. $(1,10^{-5})$-DP.
Model OPT-125M OPT-350M OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B OPT-66B
$\varepsilon=0$ $9.7$ $12.50$ $27.20$ $29.89$ $36.48$ $46.23$ $46.53$ $48.13$
$\varepsilon=1$ ${43.74_{1.49}}$ $50.69_{0.92}$ $74.41_{0.81}$ $79.30_{0.93}$ $79.53_{0.65}$ $81.35_{0.38}$ $82.03_{0.90}$ $83.87_{0.99}$
$\varepsilon=\infty$ $48.7$ $54.05$ $76.15$ $81.47$ $82.78$ $84.73$ $85.48$ $85.73$